1.         California Motors can sell preferred stock for $60 with an estimated flotation cost of $6. It is anticipated that the preferred stock will pay $5 per share in dividends.  Compute the cost of  preferred stock for the company.

2. Use the followings data for both Problem 2 and Problem 3. Power Cable Company wants you to calculate its cost of common stock. During the next 12 months, the company will  pay dividends (D1) of $3.50 per share, and the current price of  its common stock is $75 per share. The expected growth rate is 7 percent.  Flotation costs are $2.  Compute the cost of retained earnings (Ke).

3.         Use the data in problem 2 to compute the cost of new common stock.

  

4.         Sullivan Cement Company can issue debt yielding 15%. The company is paying a 33% tax rate.   What is the after-tax cost of debt?

5.    Superior Company has a $1,000 par value bond outstanding with 20 years to maturity. The bond carries an annual interest payment of $74 and is currently selling for $925. Compute the yield to maturity.
Answer #1

Question 1

Current stock price = $60

Estimated flotation cost = $6

Dividend = $5 per share

To compute the cost of preferred stock for the company:

Current stock price - Estimated flotation cost = Dividend per share / Cost of preferred stock

60 - 6 = 5 / Cost of preferred stock

54 = 5 / Cost of preferred stock

Cost of preferred stock = 5 / 54 = 0.0926 or 9.26%

Therefore cost of preferred stock for the company will be 9.26%

Note = As per Chegg policy I am only allowed to answer one question at a time. Request you to kindly place the other question separately.